Zero Turbulence Airline provides air transportation services between Los Angeles, California; and Kona, Hawaii. A single Los Angeles to Kona round-trip flight has the following operating statistics: Fuel $10,702 Flight crew salaries 8,197 Airplane depreciation 3,871 Variable cost per passenger—business class 55 Variable cost per passenger—economy class 45 Round-trip ticket price—business class 515 Round-trip ticket price—economy class 275 It is assumed that the fuel, crew salaries, and airplane depreciation are fixed, regardless of the number of seats sold for the round-trip flight. If required round the answers to nearest whole number.

a. Compute the break-even number of seats sold on a single round-trip flight for the overall product, E. Assume that the overall product is 10% business class and 90% economy class seats. Total number of seats at break-even seats

b. How many business class and economy class seats would be sold at the break-even point? Business class seats at break-even seats Economy class seats at break-even seats
Answer #1

Fixed cost = 10702+8197+3871= 22770

Weighted average contribution margin per unit = (55-45*90%+515-275*10%) = 33

a) Break even unit = 22770/33 = 690 Seats

b) Economy Class = 690*90% = 621 seats

Business class = 690*10% = 69 seats